Nothing much to report, plant bulked some more, checked trichomes and it is not done yet, mostly clear some cloudy.
Did not find any more nanners.
Values are average of the day.
DATE - EC(us/cm)
20250726 2005
20250727 2057
20250728 2201
20250729 2142
20250730 2123
20250731 2044
20250801 2061
DATE - PH
20250726 6.06
20250727 6.05
20250728 5.83
20250729 5.84
20250730 5.78
20250731 5.92
20250801 5.91
DATE - ORP (mV)
20250726 -84
20250727 -52
20250728 -50
20250729 -70
20250730 -50
20250731 -249
20250801 -269
DATE - °C - RH% (Tent Temp/RH)
20250726 24.2 62
20250727 24.0 57
20250728 23.2 58
20250729 22.6 58
20250730 23.0 53
20250731 23.2 58
20250801 25.2 59
DATE - °C (Reservoir)
20250726 21.8
20250727 21.8
20250728 21.4
20250729 21.6
20250730 22.0
20250731 22.1
20250801 22.7
DATE - CF
20250726 20.07
20250727 20.59
20250728 22.05
20250729 21.46
20250730 21.26
20250731 20.46
20250801 20.64